# Size-structured populations with growth fluctuations: Feynman–Kac formula and decoupling

### v1

@@ 8 Discussion
Our results have potential applications and extensions: First, they could be used to predict population distributions in liquid culture based on lineage averages. Second, they can help understand how size-biased sampling affects statistical analysis of single-cell data. Third, from a mathematical perspective, they introduce novel questions in stochastic processes about generalizations of the Feynman–Kac formula for non-diffusion processes and their potential application, e.g., in importance sampling.
Finally, we believe that the operator formulation of the steady-state dynamics (resulting in equations (60) and (63)) is of interest outside of the decoupling context. This could potentially be a starting point for constructing spectral representations of the population and lineage distributions, which contain information about the transient dynamics. It is also interesting to carry out a more rigorous analysis, which provides precise conditions under which these equations have unique solutions, since our Theorem has assumed the existence of such solutions.
@@ Acknowledgments
We thank Tom Chou for helpful discussions. We also thank the organizers of the 2025 Gordon Research Conference on Stochastic Physics in Biology, which helped facilitate progress on this paper.
@@ Appendix A Cell size distribution
$w_{\rm l}^{1}(y)=\int^{y}_{-\infty}dy_{b}\omega^{b}_{l}(y_{b})-\int_{-\infty}^{y}dy_{d}\omega_{l}^{d}(y_{d})$ (73)
@@
$\omega_{l}^{d}(y)=\int_{-\infty}^{y}dy_{b}f(y|y_{b})\omega_{l}^{b}(y_{b})$ (74)
In the case of symmetric division, where $u(y_{b}|y_{c})=\delta(y_{b}-y_{c}+\ln(2))$, it follows from equation (23) and equation (74), that $\omega_{l}^{d}(y)=\omega_{l}^{b}(y-\ln(2))$. Consequently, the lineage size distribution can be simplified even further as
$w_{\rm l}^{1}(y)=\int_{y-\ln(2)}^{y}dy_{b}\omega_{l}^{b}(y_{b})=F_{b}(y)-F_{b}(y-\ln(2)),$ (75)
@@
For symmetric division, we argued that there is a proportionality $w_{\rm p}(y_{c},y_{b})\propto w_{\rm l}(y_{c},y_{b})e^{y_{c}}$, hence the population cell size distribution is

### v2

@@ 8 Discussion
Our results have potential applications and extensions. First, they could be used to predict population distributions in liquid culture based on lineage averages. It is often assumed that these two types of experiments are equivalent, but rigorous comparison requires an understanding of how to link single–cell observables to population dynamics. Our results lay out a framework for doing this under certain constraints.
A second direction for future work is to explore how the decoupling and approximate–decoupling regimes we have characterized could be used to accelerate Feynman–Kac–based simulation algorithms, such as the fixed–budget Gillespie scheme of [12], for example by using the decoupled or approximately decoupled dynamics as low–variance proposals or control variates to reduce the variance of Monte Carlo estimators in these methods.
We also believe that the operator formulation of the steady-state dynamics (resulting in equations (59) and (62)) is of interest outside the decoupling context. This could potentially be a starting point for constructing spectral representations of the population and lineage distributions, which contain information about the transient dynamics.
@@ Acknowledgments
We thank Tom Chou for helpful discussions and two anonymous referees for their detailed comments, which significantly improved the quality of this manuscript. We also thank the organizers of the 2025 Gordon Research Conference on Stochastic Physics in Biology, which helped facilitate progress on this paper.
@@ Appendix A Cell size distribution
$w_{\ell}^{1}(y)=\int^{y}_{-\infty}dy_{b}w^{b}_{l}(y_{b})-\int_{-\infty}^{y}dy_{d}w_{l}^{d}(y_{d})$ (73)
@@
$w_{l}^{d}(y)=\int_{-\infty}^{y}dy_{b}f(y|y_{b})w_{l}^{b}(y_{b})$ (74)
In the case of symmetric division, where $u(y_{b}|y_{c})=\delta(y_{b}-y_{c}+\ln(2))$, it follows from equation (24) and equation (74), that $w_{l}^{d}(y)=w_{l}^{b}(y-\ln(2))$. Consequently, the lineage size distribution can be simplified even further as
$w_{\ell}^{1}(y)=\int_{y-\ln(2)}^{y}dy_{b}w_{l}^{b}(y_{b})=F_{b}(y)-F_{b}(y-\ln(2)),$ (75)
@@
For symmetric division, we argued that there is a proportionality $w_{\rm p}(y_{c},y_{b})\propto w_{\ell}(y_{c},y_{b})e^{-y_{c}}$, hence the population cell size distribution is